OpenStudy (anonymous):

What is the equation of the line passing through (-3, -9) and parallel to the line y = x + 6 in standard form? (1 points) Question 17 options: 1) x + y = -6 2) x + y = 6 3) x - y = -6 4) x - y = 6

OpenStudy (anonymous):

x-y = 6

OpenStudy (anonymous):

your a savage

OpenStudy (anonymous):

this as the same slope as y = x + 6 because it can be written y = x - 6 and it fits x = -3 y=-9

OpenStudy (anonymous):

oh - is that a compliment?

OpenStudy (anonymous):

:) can you help me with 2 more?What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x - 3 in slope-intercept form? (1 points) Question 18 options: 1) y = x + 2 2) y = x + 2 3) y = 4x + 2 4) y = -4x + 3

OpenStudy (anonymous):

this will have a slope of - 1/4 if its perpendicular so none of the above should it be parallel to ?

OpenStudy (anonymous):

if parallel then answer is 3)